STUDIES OF COLLISIONAL PROCESSES BY MICROWAVE-OPTICAL DOUBLE RESONANCE M. BARNETT and \underline{D. A. RAMSAY}, Steacie Institute of Molecular Sciences, National Research Council of Canada, Ottawa K1A 0R6.